OpenStudy (anonymous):

Two lenses of focal length +20 cm and -20cm are separated by a distance of 10 cm. Find the nature of the combines lenses.

OpenStudy (anonymous):

Will this combination behave like a convex lens and give erect but diminished image??

OpenStudy (jamesj):

The best way to answer this question is to draw an picture and trace the rays. Have you done that?

OpenStudy (anonymous):

Not really. Actually I do not how to draw for combinations....

OpenStudy (jamesj):

The other thing is this: what is the formula for the combined focal length of two lenses in contact of focal lengths f1 and f2?

OpenStudy (jamesj):

It is \[ \frac{1}{f} = \frac{1}{f_1} + \frac{1}{f_2} \] Look familiar? Now you need to modify this equation for two lenses separated a distance d. You must have this somewhere in your lecture notes or text book.

OpenStudy (jamesj):

You might also find these examples helpful: http://utopia.cord.org/step_online/st1-3/st13eiii4.htm

OpenStudy (anonymous):

Thanks for the help James. Actually it is beyond the scope of what we have been taught, so I do not have anything to go by......that is why I put it out here...☺